### v1

@@ 4.2.3 Neural Inverse Operator
We now generalize DeepSets from finite to infinite dimensions. Let $X$ be an abstract metric or Banach space and $\mu\in\mathscr{P}(X)$ be a probability measure over $X$. Consider our usual output function space $\mathcal{U}=\mathcal{U}(\Omega;\mathbb{R}^{d_{\mathrm{p}}})$. In IBVPs, $X$ is typically a space of functions defined on the boundary $\partial\Omega$. Let $\mathcal{H}$ be a latent space. Choosing operators $\Psi^{(1)}\colon X\to\mathcal{H}$ and $\Psi^{(2)}\colon\mathcal{H}\to\mathcal{U}$ leads to an operator $\Psi\colon\mathscr{P}(X)\to\mathcal{U}$ of the form (4.17). An abstract NIO [212] makes the more constrained choice
@@
Recent work has also proposed measure-to-function architectures by combining standard DeepSets with DeepONets [223, 187]. However, the NIO differs from these models because the input measures considered in [223, 187] are supported on finite-dimensional spaces. In contrast, the NIO (4.18) is the first operator learning architecture that accepts probability measures supported on functions spaces as input. Handling two levels of infinite-dimensionality, the first from $X$ and the second from $\mathscr{P}(X)$, is a notable achievement. In principle, the transformer architecture could also be lifted to infinite-dimensional state spaces and used to solve IBVPs for PDEs, but so far it has been restricted to finite dimensions due to the quadratic complexity of the attention mechanism. Regardless, all such architectures lack strong theoretical guarantees aside from some universal approximation results.
@@ 4.3 Theoretical results
The theoretical analysis of end-to-end operator learning for approximating inverse problem solution maps remains a relatively new field. Although there are scattered results for specific nonlinear inverse problems, even the theory for linear inverse problems is far from complete. A key challenge lies in the fact that input data may not necessarily belong to the domain of the inverse map being studied (see Section 4.1.2). How to reconcile this issue with the training and optimization procedures used in learning remains an open question.
@@ 4.3.1 Neural operator approximation theory: A case study on EIT
Based on results for the encoder-decoder framework [178, 228], an approximation theorem for the inverse map $\Psi^{\star}$, viewed as a mapping from a subset of $\mathcal{H}_{1}$ into $\mathcal{H}_{2}$, holds [227, Thm. 4.4, p. 786]. This theorem asserts that for every $\varepsilon>0$, there exists a generalized DeepONet $\Psi\colon\mathcal{H}_{1}\to\mathcal{H}_{2}$ such that
@@
To this end, we turn our attention to a quantitative extension approach that does use detailed stability information about the inverse map [211]. To simplify the presentation, we restrict to a setting in which $d=2$ and $\Omega\subset\mathbb{R}^{2}$ is the open unit disk. Thus, we identify the boundary $\partial\Omega$ with the unit torus $\mathbb{T}$. This aligns with Fig. 2 from Example 1. Next, we consider an infinite-dimensional set $\Gamma\subset L^{\infty}(\Omega)$ of conductivities identically equal to one outside of a compact subset of $\Omega$, uniformly bounded above and below pointwise a.e., and that have uniformly bounded $L^{1}(\Omega)$ and TV norms. This set is compact in $L^{2}(\Omega)$ and contains discontinuous conductivities of practical relevance [211, Sec. 4]. The compactness of $\Gamma$ enables us to work with uniformly accurate approximations, which interact with stability more favorably than average error estimates such as (4.22).
For any $\gamma\in\Gamma$, one can show that the NtD map $\Lambda_{\gamma}\in\mathrm{HS}(L^{2}(\mathbb{T});L^{2}(\mathbb{T}))$ is a Hilbert–Schmidt operator on $L^{2}(\mathbb{T})$ and hence a kernel integral operator. This implies that we can identify $\Lambda_{\gamma}$ with its integral kernel function $\kappa_{\gamma}\in L^{2}(\mathbb{T}\times\mathbb{T})$. Writing $\operatorname{Dom}(\Psi^{\star})\coloneqq\{\kappa_{\gamma}\,|\,\gamma\in\Gamma\}$, we seek an approximation to the inverse map $\Psi^{\star}\colon\operatorname{Dom}(\Psi^{\star})\subseteq L^{2}(\mathbb{T}\times\mathbb{T})\to L^{2}(\Omega)$ defined by $\Psi^{\star}(\kappa_{\gamma})\coloneqq\gamma$. By the Benyamin–Lindenstrauss extension theorem [229, Theorem 1.12, p. 18], the fact that $\Psi^{\star}$ maps between two Hilbert spaces and (2.3) allow us to deduce the existence of a continuous extension with log stability outside of $\operatorname{Dom}(\Psi^{\star})$ [211, Sec. 3]; recall Figs. 3 and 4. Specifically, there exists $C>0$, $\alpha>0$, and $\widetilde{\Psi}^{\star}\colon L^{2}(\mathbb{T}\times\mathbb{T})\to L^{2}(\Omega)$ such that $\widetilde{\Psi}^{\star}$ extends $\Psi^{\star}$ and
@@
The stability estimate (4.23) shows that $\widetilde{\Psi}^{\star}$ is a continuous function-to-function map and hence is amenable to universal approximation theorems for neural operators (Section 3.2). We chose to work with the FNO from (3.11), (3.14), and (3.16); other architectures would also be valid. Let $\mathsf{E}_{\sigma}$ be a compact subset of $L^{2}(\mathbb{T}\times\mathbb{T})$ with the property that $\sup_{\eta\in\mathsf{E}_{\sigma}}\lVert\eta\rVert_{L^{2}(\mathbb{T}\times\mathbb{T})}\leq\sigma$. Then by following the approach (4.10) from Section 4.1.2, we can establish a noise-robust approximation result [211, Sec. 4] which states that for every $\varepsilon>0$, there exists an FNO $\Psi\colon L^{2}(\mathbb{T}\times\mathbb{T})\to L^{2}(\Omega)$ such that
@@
By approximating the EIT inverse map using a DeepONet (3.9) adapted to linear operator inputs, the following result holds [215, Thm. 3.1, p. 7]. Let $L_{M}$ be the map $L_{M}\colon\Lambda_{\gamma}\mapsto\Lambda_{\gamma}^{M}-\Lambda_{1}^{M}$. For any $\varepsilon>0$, there exists a number of electrodes $M$ and a DeepONet $\Psi$ of the form (3.9) (with $L=L_{M}$, viewed as a function of $L_{M}(\Lambda_{\gamma})$ instead of $\Lambda_{\gamma}$) such that
@@
The DeepONet $\Psi$ in (4.25) receives as input the raw measurements $\Lambda_{\gamma}^{M}\in\mathbb{R}^{M\times M}$ shifted by the background NtD map $\Lambda_{1}^{M}$ corresponding to $\gamma\equiv 1$. The use of shifted NtD maps is for technical convenience but is not fundamentally necessary, as demonstrated in [211]. Regardless, these matrix inputs are treated as vectors in an $M^{2}$-dimensional Euclidean space rather than matrices in a Banach algebra of linear operators. It would be interesting to replace fully connected neural networks with ORN-Nets (4.15) in the branch net of $\Psi$, as these may be better suited for matrix-based data [196]; recall Section 4.2.
Due to Lipschitz stability arising from the finite dimensionality of $\Gamma$, it should further be possible to prove Lipschitz robustness to noisy NtD maps in (4.25) by adapting the ideas in [211]. However, this favorable robustness will degrade as the dimensionality of $\Gamma$—and hence the number of measurements MM—increases [231, Sec. 4.1, Thm. 5]. One would expect to only recover the logarithmic robustness shown in (4.24).
@@ 4.3.2 Finite-dimensionalizing nonlinear inverse problems
Although most analysis of deep learning for nonlinear inverse problems proceeds on a case-by-case basis, one notable exception is the work of Pineda and Petersen [198]. By transforming the conceptually infinite-dimensional nonlinear inverse problem into a finite-dimensional one, it is possible to reduce the setting to that of finite-dimensional unknown parameters and Lipschitz continuous inverse maps defined on a finite-dimensional data manifold. This transformation relies on ideas from [231]. The precise connection between the finite and infinite problems is explained in [198, Sec. 1.2.1 and Sec. 2]. A key point is that this procedure is valid for a broad class of nonlinear inverse problems. The main ingredients involve restricting the forward map to an a priori finite-dimensional set of admissible parameters and then composing a family of finite-rank “measurement” linear operators with the restricted forward map.
By carrying out this program, the original problem reduces to one about function approximation in finite dimensions. The following notable result holds true [198, Thm. 1.1]. Write $d_{\mathrm{m}}$ for manifold dimension and $d_{\mathrm{a}}$ for ambient dimension. Let $J\in\mathbb{N}$ and let $\psi^{\star}\colon\operatorname{Dom}(\psi^{\star})\subset\mathbb{R}^{d_{\mathrm{a}}}\to\mathbb{R}^{J}$ be a bounded Lipschitz mapping such that $\operatorname{Dom}(\psi^{\star})$ is a $d_{\mathrm{m}}$-dimensional relatively compact manifold satisfying certain technical conditions [198, Thm. 3.9]. There exists $\sigma_{0}>0$ such that the following holds. If $\sigma^{2}\leq\sigma_{0}^{2}$, then for every $\varepsilon>0$, there exists a ReLU neural network $\psi$ with $O(\varepsilon^{-d_{\mathrm{m}}})$ parameters such that
@@
The constant $C^{\star}$ depends on the Lipschitz constant $\operatorname{Lip}(\psi^{\star})$ of $\psi^{\star}$, the manifold $\operatorname{Dom}(\psi^{\star})$, and the maximal variance $\sigma_{0}^{2}$, but not on the error tolerance $\varepsilon$. The result as stated in (4.26) follows from [198, Thm. 3.9] and the first choice of triangle inequality (4.9) mentioned in Section 4.1.2; compare to (4.24), which uses the second triangle inequality (4.10). As usual, the proof involves extending $\psi^{\star}$ away from its domain and approximating the extension instead. The extra Lipschitz-type robustness of the network to Gaussian noise is due in part to its construction based on ideas from finite element approximation.
@@
We observe from (4.26) that the network approximation $\psi$ has a favorable noise-damping property in the following sense. First notice that $d_{\mathrm{m}}\sigma^{2}=(d_{\mathrm{m}}/d_{\mathrm{a}})\operatorname{\mathbb{E}}\lVert\eta\rVert^{2}_{\mathbb{R}^{d_{\mathrm{a}}}}$. A Lipschitz estimate for a general function on $\mathbb{R}^{d_{\mathrm{a}}}$ would only provide an average-case bound of order $\operatorname{\mathbb{E}}\lVert\eta\rVert^{2}_{\mathbb{R}^{d_{\mathrm{a}}}}$ for the perturbation $\eta$. Here, we instead see that the fine properties of the particular data manifold produce the dampening factor $d_{\mathrm{m}}/d_{\mathrm{a}}\leq 1$. This factor could be significantly smaller than $1$ if the manifold is low-dimensional compared to the ambient space. Although (4.26) corresponds only to a quantitative existence theorem, it is shown that a neural network approximation can actually be found through ERM (4.5) over noisy training pairs generated from the forward map [198, Thm. 4.3]. This trained network is accurate on average and further enjoys Lipschitz stability properties similar to those of the idealized, unattainable network $\psi$ from (4.26).
@@
To connect back to the motivation of inverse problem solvers, it is instructive to interpret the map $\psi^{\star}$ as a finite-dimensional representation of an underlying infinite-dimensional inverse map $\Psi^{\star}\colon\operatorname{Dom}(\Psi^{\star})\subseteq\mathcal{Y}\to\mathcal{U}$ between Banach spaces [231, Cor. 1]. Indeed, it is shown that the preceding theory applies to several nonlinear inverse problems corresponding to Darcy flow, Euler–Bernoulli beam, and nonlinear Volterra integral models [198, Sec. 2]. One limitation of the general approach put forth by [198] is that much of the infinite-dimensional structure, including instability as manifested by a generally non-Lipschitz modulus of continuity of $\Psi^{\star}$, is hidden in the finite-dimensionalization procedure and associated Lipschitz constants. Indeed, the neural network approximation is intertwined with the number of measurements $d_{\mathrm{m}}$. The estimation becomes harder as the parameter dimension $J$ grows because $\operatorname{Lip}(\psi^{\star})$—and hence the number of measurements $d_{\mathrm{m}}$—tends to infinity in this limit [231]. The network also must change each time $J$ or $d_{\mathrm{m}}$ changes, which violates the desirable discretization invariance principles of operator learning. Adding additional encoder-decoder structure to the approximation could reinstate consistency (see Section 3.2.1). Moreover, since the theory is only valid for finite-dimensional parameters, this precludes natural infinite-dimensional parameter models such as bounded transformations of Gaussian random fields [109, 57]. Overcoming such challenges while maintaining the wide applicability of [198] would be a substantial contribution to the field of data-driven inverse problems.

### v2

@@ 4.2.3 Neural Inverse Operator
We now generalize DeepSets from finite to infinite dimensions. Let $X$ be an abstract metric space or Banach space and $\mu\in\mathscr{P}(X)$ be a probability measure over $X$. Consider our usual output function space $\mathcal{U}=\mathcal{U}(\Omega;\mathbb{R}^{d_{\mathrm{p}}})$. In IBVPs, $X$ is typically a space of functions defined on the boundary $\partial\Omega$. Let $\mathcal{H}$ be a latent space. Choosing operators $\Psi^{(1)}\colon X\to\mathcal{H}$ and $\Psi^{(2)}\colon\mathcal{H}\to\mathcal{U}$ leads to an operator $\Psi\colon\mathscr{P}(X)\to\mathcal{U}$ of the form (4.17). An abstract NIO [233] makes the more constrained choice
@@
Recent work has also proposed measure-to-function architectures by combining standard DeepSets with DeepONets [246, 198, 197]. However, the NIO differs from these models because the input measures considered in [246, 198, 197] are supported on finite-dimensional spaces. In contrast, the NIO (4.18) is the first operator learning architecture that accepts probability measures supported on functions spaces as input. Handling two levels of infinite-dimensionality, the first from $X$ and the second from $\mathscr{P}(X)$, is a notable achievement. In principle, the transformer architecture could also be lifted to infinite-dimensional state spaces and used to solve IBVPs for PDEs, but so far it has been restricted to finite dimensions due to the quadratic complexity of the attention mechanism. Regardless, all such architectures lack strong theoretical guarantees aside from some universal approximation results.
@@ 4.3 Theoretical results
Theoretical analysis of end-to-end operator learning for approximating inverse problem solution maps remains relatively scarce. Although there are scattered results for specific nonlinear inverse problems, even the theory for linear inverse problems is far from complete. A key challenge lies in the fact that input data may not necessarily belong to the domain of the inverse map being studied (see Section 4.1.2). How to reconcile this issue with the training and optimization procedures used in learning remains an open question.
@@ 4.3.1 Neural operator approximation theory: A case study on EIT
Based on results for the encoder-decoder framework [188, 251], an approximation theorem for the inverse map $\Psi^{\star}$, viewed as a mapping from a subset of $\mathcal{H}_{1}$ into $\mathcal{H}_{2}$, holds [250, Thm. 4.4, p. 786]. This theorem asserts that for every $\varepsilon>0$, there exists a generalized DeepONet $\Psi\colon\mathcal{H}_{1}\to\mathcal{H}_{2}$ such that
@@
To this end, we turn our attention to a quantitative extension approach that does use detailed stability information about the inverse map [59]. To simplify the presentation, we restrict to a setting in which $d=2$ and $\Omega\subset\mathbb{R}^{2}$ is the open unit disk. Thus, we identify the boundary $\partial\Omega$ with the unit torus $\mathbb{T}$. This aligns with Fig. 2 from Example 1. Next, we consider an infinite-dimensional set $\Gamma\subset L^{\infty}(\Omega)$ of conductivities identically equal to one outside of a compact subset of $\Omega$, uniformly bounded above and below pointwise a.e., and that have uniformly bounded $L^{1}(\Omega)$ and TV norms. This set is compact in $L^{2}(\Omega)$ and contains discontinuous conductivities of practical relevance [59, Sec. 4.2]. The compactness of $\Gamma$ enables us to work with uniformly accurate approximations, which interact with stability more favorably than average error estimates such as (4.22).
For any $\gamma\in\Gamma$, one can show that the NtD map $\Lambda_{\gamma}\in\mathrm{HS}(L^{2}(\mathbb{T});L^{2}(\mathbb{T}))$ is a Hilbert–Schmidt operator on $L^{2}(\mathbb{T})$ and hence a kernel integral operator. This implies that we can identify $\Lambda_{\gamma}$ with its integral kernel function $\kappa_{\gamma}\in L^{2}(\mathbb{T}\times\mathbb{T})$. Writing $\operatorname{Dom}(\Psi^{\star})\coloneqq\{\kappa_{\gamma}\,|\,\gamma\in\Gamma\}$, we seek an approximation to the inverse map $\Psi^{\star}\colon\operatorname{Dom}(\Psi^{\star})\subseteq L^{2}(\mathbb{T}\times\mathbb{T})\to L^{2}(\Omega)$ defined by $\Psi^{\star}(\kappa_{\gamma})\coloneqq\gamma$. By the Benyamin–Lindenstrauss extension theorem [252, Theorem 1.12, p. 18], the fact that $\Psi^{\star}$ maps between two Hilbert spaces and (2.4) allow us to deduce the existence of a continuous extension with log stability outside of $\operatorname{Dom}(\Psi^{\star})$ [59, Sec. 3.4, Thm. 3.9]; recall Figs. 3 and 4. Specifically, there exists $C>0$, $\alpha>0$, and $\widetilde{\Psi}^{\star}\colon L^{2}(\mathbb{T}\times\mathbb{T})\to L^{2}(\Omega)$ such that $\widetilde{\Psi}^{\star}$ extends $\Psi^{\star}$ and
@@
The stability estimate (4.23) shows that $\widetilde{\Psi}^{\star}$ is a continuous function-to-function map and hence is amenable to universal approximation theorems for neural operators (Section 3.2). We chose to work with the FNO from (3.13), (3.16), and (3.18); other architectures would also be valid. Let $\mathsf{E}_{\sigma}$ be a compact subset of $L^{2}(\mathbb{T}\times\mathbb{T})$ with the property that $\sup_{\eta\in\mathsf{E}_{\sigma}}\lVert\eta\rVert_{L^{2}(\mathbb{T}\times\mathbb{T})}\leq\sigma$. Then by following the approach (4.10) from Section 4.1.2, we can establish a noise-robust approximation result [59, Sec. 4.4, Thm. 4.8] which states that for every $\varepsilon>0$, there exists an FNO $\Psi\colon L^{2}(\mathbb{T}\times\mathbb{T})\to L^{2}(\Omega)$ such that
@@
By approximating the EIT inverse map using a DeepONet (3.9) adapted to linear operator inputs, the following result holds [236, Thm. 3.1, p. 7]. Let $L_{M}$ be the map $L_{M}\colon\Lambda_{\gamma}\mapsto\Lambda_{\gamma}^{M}-\Lambda_{1}^{M}$. For any $\varepsilon>0$, there exists a number of electrodes $M$ and a DeepONet $\Psi$ of the form (3.9) (with $L=L_{M}$, viewed as a function of $L_{M}(\Lambda_{\gamma})$ instead of $\Lambda_{\gamma}$) such that
@@
The DeepONet $\Psi$ in (4.25) receives as input the raw measurements $\Lambda_{\gamma}^{M}\in\mathbb{R}^{M\times M}$ shifted by the background NtD map $\Lambda_{1}^{M}$ corresponding to $\gamma\equiv 1$. The use of shifted NtD maps is for technical convenience but is not fundamentally necessary, as demonstrated in [59]. Regardless, these matrix inputs are treated as vectors in an $M^{2}$-dimensional Euclidean space rather than as matrices in a Banach algebra of linear operators. It would be interesting to replace fully connected neural networks with ORN-Nets (4.15) in the branch net of $\Psi$, as these may be better suited for matrix-based data [216]; recall Section 4.2.
Due to Lipschitz stability arising from the finite dimensionality of $\Gamma$, it should further be possible to prove Lipschitz robustness to noisy NtD maps in (4.25) by adapting the ideas in [59]. However, this favorable robustness will degrade as the dimensionality of $\Gamma$—and hence the number of measurements MM—increases [254, Sec. 4.1, Thm. 5]. One would only expect to recover the logarithmic robustness shown in (4.24).
@@ 4.3.2 Finite-dimensionalizing nonlinear inverse problems
Although most analysis of deep learning for nonlinear inverse problems proceeds on a case-by-case basis, one notable exception is the work of Pineda and Petersen [218]. By transforming the conceptually infinite-dimensional nonlinear inverse problem into a finite-dimensional one, it is possible to reduce the setting to that of finite-dimensional unknown parameters and Lipschitz continuous inverse maps defined on a finite-dimensional data manifold. This transformation relies on ideas from [254]. The precise connection between the finite and infinite problems is explained in [218, Sec. 1.2.1 and Sec. 2]. A key point is that this procedure is valid for a broad class of nonlinear inverse problems. The main ingredients involve restricting the forward map to an a priori finite-dimensional set of admissible parameters and then composing a family of finite-rank “measurement” linear operators with the restricted forward map.
By carrying out this program, the original problem reduces to one about function approximation in finite dimensions. The following notable result holds true [218, Thm. 1.1]. Write $d_{\mathrm{m}}$ for manifold dimension and $d_{\mathrm{a}}$ for ambient dimension. Let $J\in\mathbb{N}$ and let $\psi^{\star}\colon\operatorname{Dom}(\psi^{\star})\subset\mathbb{R}^{d_{\mathrm{a}}}\to\mathbb{R}^{J}$ be a bounded Lipschitz mapping such that $\operatorname{Dom}(\psi^{\star})$ is a $d_{\mathrm{m}}$-dimensional relatively compact manifold satisfying certain technical conditions [218, Thm. 3.9]. There exists $\sigma_{0}>0$ such that the following holds. If $\sigma^{2}\leq\sigma_{0}^{2}$, then for every $\varepsilon>0$, there exists a ReLU neural network $\psi$ with $O(\varepsilon^{-d_{\mathrm{m}}})$ parameters such that
@@
The constant $C^{\star}$ depends on the Lipschitz constant $\operatorname{Lip}(\psi^{\star})$ of $\psi^{\star}$, the manifold $\operatorname{Dom}(\psi^{\star})$, and the maximal variance $\sigma_{0}^{2}$, but not on the error tolerance $\varepsilon$. The result as stated in (4.26) follows from [218, Thm. 3.9] and the first choice of triangle inequality (4.9) mentioned in Section 4.1.2; compare to (4.24), which uses the second triangle inequality (4.10). As usual, the proof involves extending $\psi^{\star}$ away from its domain and approximating the extension instead. The extra Lipschitz-type robustness of the network to Gaussian noise is due in part to its construction based on ideas from finite element approximation.
@@
We observe from (4.26) that the network approximation $\psi$ has a favorable noise-damping property in the following sense. First notice that $d_{\mathrm{m}}\sigma^{2}=(d_{\mathrm{m}}/d_{\mathrm{a}})\operatorname{\mathbb{E}}\lVert\eta\rVert^{2}_{\mathbb{R}^{d_{\mathrm{a}}}}$. A Lipschitz estimate for a general function on $\mathbb{R}^{d_{\mathrm{a}}}$ would only provide an average-case bound of order $\operatorname{\mathbb{E}}\lVert\eta\rVert^{2}_{\mathbb{R}^{d_{\mathrm{a}}}}$ for the perturbation $\eta$. Here, we instead see that the fine properties of the particular data manifold produce the dampening factor $d_{\mathrm{m}}/d_{\mathrm{a}}\leq 1$. This factor could be significantly smaller than $1$ if the manifold is low-dimensional compared to the ambient space. Although (4.26) corresponds only to a quantitative existence theorem, it is shown that a neural network approximation can actually be found through ERM (4.5) over noisy training pairs generated from the forward map [218, Thm. 4.3]. This trained network is accurate on average and further enjoys Lipschitz stability properties similar to those of the idealized, unattainable network $\psi$ from (4.26).
@@
To connect back to the motivation of inverse problem solvers, it is instructive to interpret the map $\psi^{\star}$ as a finite-dimensional representation of an underlying infinite-dimensional inverse map $\Psi^{\star}\colon\operatorname{Dom}(\Psi^{\star})\subset\mathcal{Y}\to\mathcal{U}$ between Banach spaces [254, Cor. 1]. Indeed, it is shown that the preceding theory applies to several nonlinear inverse problems corresponding to Darcy flow, Euler–Bernoulli beam, and nonlinear Volterra integral models [218, Sec. 2]. One limitation of the general approach put forth by [218] is that much of the infinite-dimensional structure, including instability as manifested by a generally non-Lipschitz modulus of continuity of $\Psi^{\star}$, is hidden in the finite-dimensionalization procedure and associated Lipschitz constants. Indeed, the neural network approximation is intertwined with the number of measurements $d_{\mathrm{m}}$. The estimation becomes harder as the parameter dimension $J$ grows because $\operatorname{Lip}(\psi^{\star})$—and hence the number of measurements $d_{\mathrm{m}}$—tends to infinity in this limit [254]. The network also must change each time $J$ or $d_{\mathrm{m}}$ changes, which violates the desirable discretization invariance principles of operator learning. Adding additional encoder-decoder structure to the approximation could reinstate consistency (see Section 3.2.1). Moreover, since the theory is only valid for finite-dimensional parameters, this precludes natural infinite-dimensional parameter models such as bounded transformations of Gaussian random fields [119, 61]. Overcoming such challenges while maintaining the wide applicability of [218] would be a substantial contribution to the field of data-driven inverse problems.